new york city typically has rain on about 16% of days in October what is the probability that it will rain on exactly 5 days in October and 15 days?


1
Expert's answer
2022-05-13T04:52:11-0400

The number of rainy days, "X" can be represented by a binomial distribution with "n=31" trials (the number of days in the month of October), success probability "p=0.16" (representing a rainy day) and failure probability "q=1-p=0.84" .

Therefore,


"P(5)=_{n}C_{5}p^5q^{n-5}=_{31}C_{5}(0.16)^5(0.84)^{26}\approx0.191""P(15)=_{n}C_{15}p^{15}q^{n-15}=_{31}C_{15}(0.16)^{15}(0.84)^{16}\approx0.0000213"